# LLM Personas as a Substitute for Field Experiments in Method Benchmarking

### v2

@@ B.3 Extension to heteroscedastic Gaussian reduced-form kernels
Moreover, for $L$ i.i.d. aggregate evaluations (i.e., $Q_{\mathrm{pers}}(\cdot\mid w)^{\otimes L}$),
$D_{\mathrm{KL}}\!\Big(Q_{\mathrm{pers}}(\cdot\mid w)^{\otimes L}\,\Big\|\,Q_{\mathrm{pers}}(\cdot\mid w^{\prime})^{\otimes L}\Big)=L\,D_{\mathrm{KL}}\!\big(Q_{\mathrm{pers}}(\cdot\mid w)\,\big\|\,Q_{\mathrm{pers}}(\cdot\mid w^{\prime})\big).$
@@
For the product statement, note that $P^{\otimes L}$ and $Q^{\otimes L}$ have densities $p_{L}(o_{1:L})=\prod_{i=1}^{L}p(o_{i})$ and $q_{L}(o_{1:L})=\prod_{i=1}^{L}q(o_{i})$, so
$\log\frac{p_{L}(O_{1:L})}{q_{L}(O_{1:L})}=\sum_{i=1}^{L}\log\frac{p(O_{i})}{q(O_{i})}.$
Taking expectation under $O_{1:L}\sim P^{\otimes L}$ and using independence,
$D_{\mathrm{KL}}(P^{\otimes L}\|Q^{\otimes L})=\sum_{i=1}^{L}\mathbb{E}_{O_{i}\sim P}\!\left[\log\frac{p(O_{i})}{q(O_{i})}\right]=L\,D_{\mathrm{KL}}(P\|Q).$
This proves the lemma. ∎
@@ Appendix C Experiment: Discriminability Calibration for the Persona-Based Ad Benchmark
For each artifact $w\in W_{0}$, we run $n_{0}$ independent repeats of TextBO’s evaluation call:
$o_{1}(w),\dots,o_{n_{0}}(w)\;\;\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\;\;Q_{\mathrm{pers}}(\cdot\mid w),$
where each repeat resamples a fresh panel of $n_{0}=200$ personas (all other evaluation components are held fixed, including the judge prompt and model version). We compute the empirical mean and variance:
$\widehat{\mu}(w)=\frac{1}{n_{0}}\sum_{t=1}^{n_{0}}o_{t}(w),\qquad\widehat{\sigma}^{2}(w)=\frac{1}{n_{0}-1}\sum_{t=1}^{n_{0}}(o_{t}(w)-\widehat{\mu}(w))^{2}.$
For each scenario, for all pairs $w,w^{\prime}\in W_{0}$ with $d_{W}(w,w^{\prime})\geq r$, we compute the pairwise signal-to-noise proxy:
$\widehat{\mathrm{SNR}}(w,w^{\prime})\;=\;\frac{\big(\widehat{\mu}(w)-\widehat{\mu}(w^{\prime})\big)^{2}}{\widehat{\sigma}^{2}(w)+\widehat{\sigma}^{2}(w^{\prime})}.$
There are many possible combinations of $w,w^{\prime}$. Kang and Yoganarasimhan (2025) reported results for ten steps of prompt improvement from the initial prompt; we tested ten improvement variants for each step. This totals 100 TextBO prompt improvement cases per scenario. We then define a conservative discriminability estimate as a low quantile of these values, instead of taking a literal infimum:
$\widehat{\kappa}_{Q}\;=\;\mathrm{Quantile}_{q}\Big(\{\widehat{\mathrm{SNR}}(w,w^{\prime}):\;d_{W}(w,w^{\prime})\geq r\}\Big),$
with $q=0.05$. We report $\widehat{\kappa}_{Q}$ per scenario and use the minimum across scenarios as the overall conservative estimate. The predicted number of independent evaluation calls per prompt required to decide between two $r$-separated artifacts is
$n_{\mathrm{req}}\;=\;\left\lceil\frac{2}{\widehat{\kappa}_{Q}}\log\frac{1}{\delta}\right\rceil.$
Table 1 provides $\widehat{\kappa}_{Q}$ and $n_{\mathrm{req}}$ for each scenario. Across the scenarios, $n_{\text{req}}$ varies from 20 to 513, with an average of 204.38. This shows that (Kang and Yoganarasimhan, 2025)’s choice of random sampling 200 personas for training was a legitimate choice.
Table 1: $\widehat{\kappa}_{Q}$ and Sample Size for $\delta=0.05$
@@
$\widehat{\kappa}_{Q}$ 0.00508 0.02320 0.13178 0.00942 0.02955 0.13069 0.0046 0.07548 0.051225
$n_{\text{req}}$ 513 113 20 280 89 20 565 35 204.375

### v3

@@ Appendix C Experiment: Discriminability Calibration for the Persona-Based Ad Benchmark
There are many ways of choosing the sampling distribution of $(w,w^{\prime})$ pairs. Kang and Yoganarasimhan (2025) reported results for ten steps of prompt improvement from the initial prompt; we tested ten improvement variants for each step. This totals 100 TextBO prompt improvement cases per scenario. We then estimate $\widehat{\kappa}_{Q}(q)$ per scenario. By Lemma 5.2, the predicted number of independent persona samples per prompt required to decide between two $r$-separated artifacts is
$L_{\mathrm{req}}\;=\;\left\lceil\frac{2}{\widehat{\kappa}_{Q}(q)}\log\frac{1}{\delta}\right\rceil.$
Table 1 provides $\widehat{\kappa}_{Q}(q)$ and $L_{\text{req}}$ for each scenario. Across the scenarios, $L_{\text{req}}$ varies from 46 to 1180, with an average of 460.25. This shows that the persona sample size choice of 200 is not too bad, but increasing it to 500 could have been a more conservative choice.
Table 1: $\widehat{\kappa}_{Q}(q)$ and Sample Size for $q=0.05$ and $\delta=0.05$
@@
$\widehat{\kappa}_{Q}(q)$ 0.00508 0.02320 0.13178 0.00942 0.02955 0.13069 0.0046 0.07548 0.051225
$L_{\text{req}}$ 1180 259 46 637 203 46 1303 80 469.25
Appendix D Further discussions
D.1 Further discussions on contribution
Is AO+MB equivalence close to tautological?
The paper’s main contribution is not that the factorization exists once you assume it, but that the “drop-in substitute” claim can be identified exactly, and only, from two protocol-level benchmark-hygiene conditions that are concrete and auditable: what information is exposed to methods (AO) and whether the evaluator is provenance/identity independent (MB). The equivalence is valuable precisely because it collapses an otherwise ambiguous debate (“do personas ‘match’ humans?”) into two checkable failure modes that are independent in practice and routinely violated in real evaluation pipelines. In particular, MB is not a cosmetic assumption: it formalizes that the benchmark is even a well-defined environment, and the paper explicitly discusses why MB is fragile in human/LLM-judge settings unless enforced by design.
Moreover, the result is not “MB alone implies JPC.” The necessity of AO is nontrivial: even if two pipelines induce the same aggregate score distribution, leaking micro-level information can allow adaptive methods to distinguish and exploit differences. The paper includes an explicit counterexample demonstrating that violating AO can break “just panel change” even when the aggregate kernels match (Proposition B.1 in Appendix B.1). That is exactly the kind of conceptual point that is easy to miss without a formal interface-level definition and a necessity proof.
D.2 Further discussions on extensions
Would more empirical validation strengthen the arguments in this paper?
The central claim in this paper is an identification theorem about benchmark interfaces, not an empirical claim that a particular persona pipeline tracks humans. The paper includes a proof-of-concept $\mathrm{\kappa Q}$ calibration (Appendix C) to demonstrate the intended workflow: $\mathrm{\kappa Q}$ is designed to be measured and to turn ”persona quality” into a budget/design question. Requiring extensive multi-domain experimentation is therefore more about strengthening the paper’s applied guidance than about validating the correctness of the main theorem.
Can sample complexity easily be extended beyond pairwise?
Pairwise comparison is the primitive building block, and this is easily extended beyond pairwise comparison. For example, If if the workflow is ”compare $K$ fixed artifacts,” pairwise error control can be achieved with a union bound by allocating per-comparison failure probability $\delta^{\prime}=\delta/C(K,2)$) (or $\delta/\mathrm{K}$ for tournament-style brackets), which yields only a log-factor change in required samples.